### v1

@@ 5 UN–DID with Staggered Adoption
The above process is repeated for all relevant $2\times 2$ group–time cells. These ATT(g,t) cells will then be aggregated together to get an overall estimate of the ATT based on Equation (60). $w_{g,t}$ are the weights associated with each of the ATT(g,t) cells. Common weighting schemes include equal weighting for each cell and population-based weighting. Refer to Callaway and Sant’Anna (2021) and Xiong et al. (2023) for details.
@@ 6 Empirical Examples
6.1 The Impact of Marijuana Legalization on Body Mass Index
We revisit the question asked in Sabia et al. (2017) to see how comparable the UN–DID estimates of an ATT may be in comparison to conventional methods. Specifically, we ask whether whether legalizing marijuana has any impact on body mass index (BMI).
We use data from the Behavioral Risk Factor Surveillance System. This repeated cross-section has records from people of all ages. However, we restrict our attention to those aged 25 or older. We use data from 2006–2013 from New York and the District of Columbia (DC). DC relaxed their marijuana laws in 2010. Thus we have a $2\times 8$ design with an equal number of pre-treatment and post-treatment years. There are 84,783 observations in the final sample.
With this data, we estimate four different treatment effects: from a conventional DID regression and with UN–DID, with and without controls. The control variables are female and categorical variables for age, race, and education. The conventional DID model we estimate is shown in Equation (61). We estimate the UN–DID ATT using the alternate specification given in equations (53) and (53).
@@ 6.1 The Impact of Marijuana Legalization on Body Mass Index
Table 1 presents the estimates. As expected, without covariates, we get exactly the same coefficients from both the conventional and UN–DID estimates, with an estimated ATT of 0.2560 and a standard error of 0.0853. This suggests that there was a statistically significant increase in BMI after the legalization of marijuana, in contrast to the findings in Sabia et al. (2017).
When we add covariates, the estimated coefficient drops considerably, to 0.1322 for the conventional DID and to 0.1204 for UN–DID. Again, this difference between the two estimates is expected, especially because there are 25 covariates in the model. The standard errors are also slightly smaller, at 0.0807 for the conventional model and 0.0811 for UN–DID. The p-value for the conventional DID is just above the 10% threshold, at 0.102, suggesting marijuana legalization did not have a statistically significant impact on BMI.
Table 1: Effect of Marijuana Legalization on BMI
Conventional UN–DID Conventional UN–DID
@@ 6.2 The Impact of Merit Scholarships on College Attendance
To highlight the use of the new method proposed, we revisit the analysis of “merit scholarships.” These scholarships vary in their design, generally awarding scholarships to high achieving high school students to enroll in a college within their home state. See Deming and Dynarski (2010) and the references therein for a nice survey. We use the dataset analyzed as an empirical example in Conley and Taber (2011), which estimates the average impact of 10 different merit scholarship programs with different adoption dates.
Specifically, we estimate how the likelihood of an individual being a college graduate changes as a result of their home state adopting a merit scholarship program. We first estimate this effect individually for each of the 10 treated states. This allows us to compare 10 distinct estimates of treatment effects using both the UN–DID procedure and the the conventional procedure. There are 51 “states” (including DC) in the dataset and 12 years of data from 1989 to 2000. Across all states and years, there are 42,161 individual-level observations, representing a repeated cross-section.
Ten states in the sample adopt a merit scholarship, with adoption dates that vary greatly. Four states adopt the scholarship in 1991, 1993, 1996, and 1999, respectively, while two states each adopt the scholarship in years 1997, 1998, and 2000. For each treated state, a control state was chosen from the set of 41 untreated states. The control states were chosen to roughly match the number of treated and untreated observations in each model. We use only one treated state to highlight the properties of the UN–DID estimator rather than to estimate the best possible counterfactuals. We estimate the treatment effects both with and without covariates. Specifically, using only data from treated state $T$ and untreated state $U$, we estimate the model:
@@
The coefficients $\delta_{1},\delta_{2}$, and $\delta_{3}$ are estimated only when the covariates are included. The remaining variables are defined the standard way: $treat_{s}$ is set equal to 1 for observations in state $T$ and is set equal to 0 for observations in state $U$. Similarly, $post_{t}$ is set equal to 1 for years greater than and equal to the year when state $T$ first offered a merit scholarship. Finally, $did_{st}$ is the product of $treat_{s}$ and $post_{t}$. We estimate all these treatment effects, or $\widehat{did}_{st}$, and their standard errors using both the pooled dataset and the UN–DID estimates and standard errors by artificially siloing the data.
@@
The results of these estimates are found in Table 2. Panel A shows the estimates for the models without covariates. The first thing to notice is that the treatment effects are identical when we estimate using either the conventional method with the pooled data or the UN–DID method with the unpooled data. Additionally, the standard errors always agree to the third decimal place, and agree to the fourth decimal place in all but one of the estimates. These results match the simulation results presented in Section 3.3.
@@
Panel B of Table 2 shows the estimates with covariates included. Here, some of the treatment effects are estimated to be very similar using either the conventional pooled approach or the UN–DID approach. For instance, for the treatment of state 88 versus state 11, the conventional estimate is 0.1811 and the UN–DID estimate is 0.1819. However, some of the estimates are not as close. The treatment effect of state 61 versus state 62 is estimated to be 0.0370 with the conventional approach and 0.0314 with the UN–DID approach. The estimates for the standard errors differ considerably more. These findings are consistent with the results in Section 4.4.
Additionally, we attempt to estimate a single treatment effect using data from all years and all groups. We do this using three methods, two with pooled data and one with unpooled data. The first pooled method estimates a two-way fixed effect (TWFE) difference-in-differences model, similar to Equation 62 but replacing the $treat_{s}$ and $post_{t}$ terms with state and year fixed effects. This replicates the original analysis in Conley and Taber (2011). Given the staggered adoption, the TWFE estimates contain “forbidden comparisons.” Accordingly, we also estimate the ATT using the CSDID command, which implements the Callaway and Sant’Anna (2021) procedure. Finally, we estimate the same ATT, using only the non-forbidden comparisons estimated with the UN–DID method by again pretending that each state’s data is siloed.
The aggregate ATT estimates are shown in Panel C of Table 2. With no covariates we find that all three estimates are quite similar. Specifically, the TWFE estimate is 0.0374, the CSDID estimate is 0.0339, and the UN–DID estimate is 0.0361. We do not expect the CSDID and UN–DID estimates to entirely agree, for two reasons. The first is that the CSDID estimate is doubly robust and the UN–DID estimate is not. The second is the “treatment unit” in CSDID is the year of adoption, whereas we regard the state as the “treatment unit” because that is the level of siloing of the data. With covariates, we see that the TWFE and CSDID estimates are very similar, at 0.0337 and 0.0339, respectively. The UN–DID estimate is slightly smaller at 0.0295. One reason for this difference is that the UN–DID approach in this case involves 51 separate regressions estimating 153 coefficients for the controls, whereas the pooled approaches need to estimate only three coefficients to handle the controls. Taken together, these estimates suggest that the UN–DID approach with unpooled data gives very similar estimates to conventional methods using pooled data.
@@
Treated Control Year Conventional UN–DID Conventional SE UN–DID SE
@@
Treated Control Year Conventional UN–DID Conventional SE UN–DID SE
@@
Years Covariates UN–DID TWFE CSDID
All No 0.0361 0.0374 0.0339
All Yes 0.0295 0.0337 0.0339
@@ 7 Conclusion
In this paper, we propose an extension to the conventional regression-based method for difference-in-differences to settings with unpoolable data called UN–DID. Using simulations and analytical proofs, we show that the proposed method results in identical coefficients and standard errors to the conventional method with no covariates. We also show that it is unbiased. When time-invariant covariates are added, the coefficients can be the same as the conventional ones, but the standard errors differ. Time-varying covariates cause neither the coefficients nor the standard errors to be equal with small sample sizes but converge as sample size increases. We extend the UN–DID method to a setting with staggered adoption and verify the results from the simulation study and analytical proofs with an empirical example.

### v2

@@ 5 UN-DID with Staggered Adoption
The above process is repeated for all relevant $2\times 2$ group–time cells. These ATT(g,t) cells are then aggregated together to get an overall estimate of the ATT based on Equation (54). $w_{g,t}$ are the weights associated with each of the ATT(g,t) cells.
@@
Common weighting schemes include equal weighting for each cell and population-based weighting. Refer to Callaway and Sant’Anna (2021) and Xiong et al. (2023) for details.
6 Multiple Silos
Up until Section 5, we have concentrated on settings in which there are only two silos. However, even with common timing it is typically the case that many treated and many comparison silos will be used. In that case, it is easier to estimate the average treatment effect by means of a second stage regression. The details differ slightly depending on whether there is common or staggered adoption.
6.1 Common Adoption
Imagine that there are two treated silos, A and B, and one control silo C. For these purposes, consider that the various silos contain the individual level observations, and that the data sharing agreements permit summary statistics to be extracted, but no new information can be added to a silo’s secure environment. In that case, calculation of the treatment effect must occur on a fourth computer, call this a server. Let us also assume that we are not interested in estimating dynamic treatment effects, so that aggregating time points into pre- and post-intervention will suffice. In that case, we need three pieces of information from each silo: the difference between the pre-period and the post-period, as estimated in Equation (42), the standard error for this difference, and possibly the weight for the silo. We typically recommend the weight for silos $s$ of $w_{s}=N^{post}_{s}/(N^{post}_{s}+N^{pre}_{s})$.
The first stage of the analysis creates a dataset which contains this information, along with treatment status for each of the three silos. In this case, the dataset would have three observations, and four variables: $\textrm{diff}_{s}$ is the within silo pre-post difference, potentially after controlling for covariates; $se_{s}$ is the standard error for $\textrm{diff}_{s}$; $w_{s}$ is the silo weight; and $d_{s}$ is whether that silo is treated $d_{s}=1$ or not $d_{s}=0$. We can calculate the ATT by estimating the regression:
$\textrm{diff}_{s}=\alpha+\beta d_{s}+\epsilon_{s}.$ (55)
In this regression, $\beta$ is the ATT, which will give the same estimate that one would get from a regression using poolable data when the groups are of equal size. When the groups are of unequal size, a weighted least squares version of the regression is needed to match the regression with pooled data. This second stage regression approach is particularly useful for cluster robust inference. This is being explored in a companion paper in development by the authors of this paper.
6.2 Staggered Adoption
Staggered adoption means an additional computational burden with multiple silos. Each individual ATTg,t can be calculated as before, and then an overall ATT can be estimated as a weighted average (see Section 5). To fix ideas, consider the simplest staggered adoption setting, where there are three groups (A,B,C) and three time periods (1,2,3). Group A is first treated in period 2 ($g=2$). Group B is first treated in period 3 ($g=2$). Group C is never treated ($g=\infty$). In this 3$\times$3 example, without covariates, and using the never-treated cells as controls, the three $ATT_{gt}$s are:
$\begin{array}[]{rcl}ATT_{22}&=&(\bar{Y}_{A2}-\bar{Y}_{A1})-(\bar{Y}_{C2}-\bar{Y}_{C1})\\[8.0pt] ATT_{23}&=&(\bar{Y}_{A3}-\bar{Y}_{A1})-(\bar{Y}_{C3}-\bar{Y}_{C1})\\[8.0pt] ATT_{33}&=&(\bar{Y}_{B3}-\bar{Y}_{B2})-(\bar{Y}_{C3}-\bar{Y}_{C2})\end{array}$
In this example, we need to calculate differences in mean outcomes within each silo for different sets of years for each silo. Specifically, we need to estimate the difference between year 2 and 1, and 3 and 1 for silo A. For silo B, we only need the difference between year 3 and 2. Finally, for silo C, we need all three of these differences. Similar to the multiple silo with common timing example, we propose storing these differences, and their standard errors, in a separate matrix. The staggered treatment timing requires us to also keep track of the $g$ and $t$ for each difference that we calculate, as well as whether a given silo is a treatment or control silo. To fill this matrix, each of the within-silo, two-year differences can be calculated using a separate regression. Table 1 illustrates this matrix for a simple 3$\times$3 example.
Table 1: Second Stage Matrix
Contrast Silo d G T diff
A22 A 1 2 2 $\bar{Y}_{A2}-\bar{Y}_{A1}$
A23 A 1 2 3 $\bar{Y}_{A3}-\bar{Y}_{A1}$
B33 B 1 3 3 $\bar{Y}_{B3}-\bar{Y}_{B2}$
C22 C 0 2 2 $\bar{Y}_{C2}-\bar{Y}_{C1}$
C23 C 0 2 3 $\bar{Y}_{C3}-\bar{Y}_{C1}$
C33 C 0 3 3 $\bar{Y}_{C3}-\bar{Y}_{C2}$
The various ATTg,t terms, can then be estimated using this matrix estimating this regression:
$\textrm{diff}_{s,g,t}=\alpha+\beta d_{s,g,t}+\epsilon_{s,g,t}\textrm{ if }G=g\textrm{ and }T=t.$
In this regression $\beta$ estimates the ATTg,t given by the sample restriction. For the 3$\times$3 example, the three restrictions needed would be G=2, T=2; G=2, T=3; and G=3, T=3. We have developed a software package, UNDID which implements this algorithm. The package was built in Julia, but also has implementations in R and Stata, the package is available here: https://github.com/ebjamieson97/Undid.jl.
@@ 7 Empirical Examples
We now consider two empirical examples, one with common adoption timing and one with staggered adoption. We include these examples to highlight the properties of the UN-DID and conventional DID methods. In contrast to our Monte Carlo simulations, we do not know the true effects of the policies we analyze here. Using survey data that is pooled across US states, we can estimate the ATT using both conventional DID and UN-DID (treating the data as if it were siloed by state), and then compare the ATTs and SEs estimated using these two approaches. Our purpose here is to compare the two methods with real-world data - not to answer the substantive question. Accordingly, we diverge somewhat from the designs in the original analyses, and our analysis falls well short of complete empirical analyses of the substantive treatment effects.
7.1 The Impact of Medical Marijuana Legalization on Body Mass Index
We revisit the question asked in Sabia et al. (2017) of how access to medical marijuana impacts body mass index. The original study investigates the role of medical marijuana laws in 21 US states and finds that access to medical marijuana decreases the likelihood of being obsese. To assess the performance of UN-DID in a 2$\times$×T setting with common treatment timing, we explore the impact of the District of Columbia’s (DC) medical marijuana law, using New York State as the control.
Data come from the Behavioral Risk Factor Surveillance System (BRFSS). This repeated cross-sectional survey is pooled across US states but can also be analyzed as if the data were siloed by state. We use data from 2006–2013 from New York and DC, which enacted and enforced their medical marijuana law in 2010. We therefore have a $2\times 8$ design with an equal number of pre-treatment and post-treatment years. We restrict our attention to respondents aged 25 or older, resulting in a final sample of 84,783 observations.
With these data, we estimate treatment effects using four different models: a conventional DID regression and UN-DID, each with and without controls. The control variables are female and categorical variables for age, race, and education, which are time-varying covariates in a repeated cross-sectional dataset. The UN-DID estimates are calculated by artificially siloing the data by state. The conventional DID model is shown in Equation (56).
@@ 7.1 The Impact of Medical Marijuana Legalization on Body Mass Index
We also the UN-DID ATT by first estimating the following model separately for New York and DC.
$\footnotesize\text{BMI}_{ist}=\beta_{1}\text{post}_{t}+\beta_{2}\text{post}_{t}+\delta_{1}\text{female}_{ist}+\delta_{2}\text{RACE}_{ist}+\delta_{3}\text{AGE}_{ist}+\delta_{4}\text{EDUC}_{ist}+\epsilon_{ist}.$ (57)
Table 2 presents the ATT and standard error estimates. As expected, without covariates the conventional DID and UN-DID yield the same estimates: an estimated ATT of 0.2560 and a standard error of 0.0853. When we add covariates, the estimated coefficient drops to 0.1322 for the conventional DID and 0.1204 for UN-DID. These estimates are neither substantively nor statistically different from one another. The small difference between them is expected given that the models include 25 covariates which are restricted to common slopes with conventional DID and allowed to vary as silo-specific slopes with UN-DID. The standard errors are also slightly different, but substantively equivalent: 0.0807 for the conventional DID and 0.0811 for UN-DID.
This empirical example demonstrates the strong performance of UN-DID relative to conventional DID in a 2$\times$×T setting with common or single treatment timing. We reiterate that our effect estimates in this simplified example should not be interpreted as responding to the substantive question. However, this analysis using real-world, artificially siloed data demonstrates that UN-DID can accurately estimate the ATT and SE as well as conventional DID, both with and without covariates.
Table 2: Effect of DC’s Medical Marijuana Legalization on BMI
Conventional DID UN-DID Conventional DID UN-DID
@@ 7.2 The Impact of Merit Scholarships on College Attendance
To highlight the use of UN-DID in a staggered adoption setting, we revisit the analysis of merit scholarships from Conley and Taber (2011). These authors estimate the average impact of 10 state-level merit scholarship programs with different adoption dates. These scholarships vary in their design, generally awarding scholarships to high-achieving high school students to enroll in a college within their home state (see Deming and Dynarski (2010) and the references therein for details).
Data come from the Current Population Survey (CPS), another repeated cross-sectional survey that is pooled across US states. There are 51 “states” (including DC) and 12 years (1989-2000) in the dataset, for a total of 42,161 individual-level observations. We estimate the ATT of a state’s adoption of a merit scholarship program on the likelihood of a student in that state graduating college. We first estimate this effect individually for each of the 10 treated states, and then combine those for an aggregate ATT. This allows us to compare 11 distinct treatment effect estimates using both UN-DID and conventional DID.
Ten states adopt a merit scholarship across varying adoption dates. Four states adopt the scholarship in 1991, 1993, 1996, and 1999, respectively, while two states each adopt in 1997, 1998, and 2000. For each treated state, one control state was chosen from the set of 41 untreated states to roughly match the number of treated and untreated observations in each model. We use only one treated state to highlight the properties of the UN-DID estimator rather than to estimate the best possible counterfactuals. We estimate the treatment effects both with and without covariates. Specifically, using only data from treated state $T$ and control state $C$, we estimate the model:
@@
The coefficients $\delta_{1},\delta_{2}$, and $\delta_{3}$ are estimated only when the covariates are included. The remaining variables are defined the standard way: $treat_{s}$ is set equal to 1 for observations in state $T$ and is set equal to 0 for observations in state $U$. Similarly, $post_{t}$ is set equal to 1 for years greater than and equal to the year when state $T$ first offered a merit scholarship. Finally, $did_{st}$ is the product of $treat_{s}$ and $post_{t}$. We estimate treatment effects, or $\widehat{did}_{st}$, and standard errors using the pooled dataset.
We then artificially silo the data to estimate treatment effects and standard errors using UN-DID. Specifically, for each treated silo we estimate
$\text{college}_{ist}=\beta_{1}\text{pre}_{st}+\beta_{2}\text{post}_{st}+\delta_{1}\text{black}_{ist}+\delta_{2}\text{asian}_{ist}+\delta_{3}\text{male}_{ist}+\epsilon_{ist}\text{ for }s\in(C,T).$ (59)
Here, $\text{pre}_{t}$ is equal to one in the years before the merit program was adopted, and equal to zero afterwards. Conversely, $\text{post}_{t}$ is equal to zero in the years before the merit program was adopted, and equal to one afterwards. We estimate the same regression for the control state used for each treated state. We then calculate the treatment effect using Equation (8) and the standard error using Equation (9).
@@
The results are found in Table 3. Panel A shows the estimates for the models without covariates. The treatment effect and standard error estimates are identical to the fourth decimal place using conventional DID with the pooled data or UN-DID with the unpooled data. These results match the simulation results presented in Section 3.3.
@@
Panel B of Table 3 shows the treatment effect estimates conditional on covariates. Here, some of the individual state treatment effect estimates are very similar using conventional DID and UN-DID. For instance, for state 88, the conventional DID effect estimate is 0.1811 and the UN-DID estimate is 0.1819. However, some of the estimates are not as close. The estimated treatment effect for state 61 is 0.0370 using conventional DID and 0.0314 with UN-DID. Even though the difference between the estimates is larger, they are not statistically different and may not be substantively different from each other. The estimates of the standard errors differ more between conventional DID and UN-DID. Again, these differences are not large enough to impact inference in this example. These findings are consistent with the results in Section 4.5.
Lastly, we estimate average treatment effects using data from all groups and all necessary years per group. We do this using two methods, both including covariates. We estimate the ATT using the CSDID command, which implements the Callaway and Sant’Anna (2021) procedure. We again pretend that each state’s data are siloed, and estimate the ATT using UN-DID. We do this using our UNDID command, which excludes forbidden comparisons. Simplifying the procedure, we estimate an ATT for each post treatment period for each treated state. This is done by estimating a model like in Equation (59) for each treated state, but using only one pre-treatment and one post-treatment year. Using those same years, we estimate the same model for each of the control states. The state-specific first differences $\beta_{2}-\beta_{1}$ are then calculated, and added to a table, like Table 1 as described in Section 5. We can then estimate each ATTg,t by running a properly specified regression like the one in Equation (55). For both CSDID and UN-DID we estimate one ATT by taking a simple average of the ATTg,t and one by averaging over the $g$ groups. Additionally, we calculate cluster-robust standard errors, assuming state level clustering, for all ATT estimates. The UN-DID standard errors use the jackknife discussed in Section 2.1.
The aggregate ATT estimates are shown in Panel C of Table 3. We find the simple average ATT estimates from the two procedures to be quite similar. The UN-DID ATT is 0.0485 and the CSDID ATT is 0.0464. The two standard errors are also very close to one another. The group-averaged ATTs are not as similar to one another, but still close. The UN-DID ATT is 0.0459, remaining close to the simple average ATT. The CSDID ATT drops to 0.0339. The standard errors are also not as similar, with the UNDID standard error equal to 0.0188 and the CSDID one equal to 0.0118. Given the UN-DID standard error is from a cluster jackknifed and the CSDID standard error is not, it makes sense that it is larger as the 51 states are unbalanced in the number of observations. This example, using 51 artificial silos, shows how accurately the UN-DID procedure can estimate an ATT.
@@
Treated Control Year DID UN-DID DID SE UN-DID SE
@@
Treated Control Year DID UN-DID DID SE UN-DID SE
@@
Agg. UN-DID ATT UN-DID SE CSDID ATT CSDID SE
group 0.0459 0.0188 0.0339 0.0118
simple 0.0485 0.0110 0.0464 0.0117
@@ 8 Conclusion
Difference-in-differences is commonly used to estimate treatment effects, particularly for policies or other interventions that vary at the jurisdictional (e.g., state, country) level. However, this approach was previously infeasible in settings where the data are siloed by jurisdiction. In this paper, we propose the unpoolable DID (UN-DID) method, assess its performance, and demonstrate its utility in settings with siloed data. We also formalize the assumption of data poolability in the DID context, which was previously implied but not directly articulated. Specifically, we show that UN-DID is an unbiased and reliable extension to the conventional regression-based DID method in settings with unpoolable data.
We began by presenting analytical proofs to demonstrate the equivalence of UN-DID and conventional DID, both with and without covariates. We then extended our assessment to simulated panel data using a series of Monte Carlo simulations (incorporating a range of DGPs and sample sizes). These experiments permitted a direct assessment of UN-DID’s performance and unbiasedness compared to both conventional DID and the true treatment effects. To better understand UN-DID’s performance with real-world data, we concluded by comparing the performance of UN-DID and conventional DID using pooled data from existing empirical examples, treating these data as unpoolable.
Across all assessments, we showed that UN-DID and conventional DID consistently return identical treatment effect estimates and standard errors in models without covariates. The ATTs remain unbiased when we include time-invariant and time-varying covariates. The SEs are also unbiased but not numerically equal, converging to the true value as the sample size increases. These findings suggest that UN-DID is a valid approach to DID when data are unpoolable.
Our work contributes to a growing body of literature developing DID methods for use in increasingly complex policy and data environments. UN-DID offers an important methodological tool for researchers assessing multi-jurisdictional (or cross-jurisdictional) policy effects. By enabling the use of DID estimation in siloed data settings, UN-DID effectively opens the door to new research questions and new evidence on policy impacts.
The UN-DID software packages in Stata, R, and Julia will facilitate the practical application of UN-DID by a range of different end-users; the accompanying user guide will further support this aim while additionally showcasing the use of UN-DID in a truly siloed setting. Taken together, we expect our development of UN-DID (both as an analytical method and an associated software package) to catalyze innovative cross-jurisdictional policy evaluation in siloed data settings, thereby eliminating a common barrier for an otherwise widely used and powerful method.